OpenStudy (anonymous):

You know that there are 46 dimes and quarters in a jar. When you count up the money you find you have a total of $9.40. How many dimes are there? How many quarters?

OpenStudy (amistre64):

setup a system of equations d + q = 46 .10d + .25q = 9.40

OpenStudy (anonymous):

So do you solve for D or Q? and how

OpenStudy (amistre64):

i would solve for d.imes and q.uarters yes. and your material should give you at least 3 ways to approach a solution. use the one you are most comfortable with.

OpenStudy (amistre64):

elimination seems to be the simplest approach to me since the top is practically bare of coeffs ... but many like subsitution better

OpenStudy (amistre64):

let d = (23+n) and q=(23-n) d + q = 46 23 + n + 23 - n = 46 ... by default substitute them into the second equation and solve for n.

OpenStudy (anonymous):

Ummm ok how do you do substitution? I know what substitution is and I know the equation, I just don't know how to use the two together.

OpenStudy (anonymous):

Where does 23 come from?

OpenStudy (amistre64):

one type of substitution i just mentioned the other type is to solve one equation for a term, and replace it in the other equation if ax + by = c, we can solve for a: a = (c-by)/x this is one type of relationship that must be satisfied so we simply let all a's be (c-by)/x

OpenStudy (amistre64):

23 is a convenient number since 23 + 23 = 46 and n-n = 0 its just useful to define them that way

OpenStudy (amistre64):

d = 23 + n + q = 23 - n ------------ 46 = 46 + 0

OpenStudy (amistre64):

.10d + .25q = 9.40 substitute for d and q .10(23+n) + .25(23-n) = 9.40 now we only have one variable to solve

OpenStudy (anonymous):

n= -9

OpenStudy (amistre64):

very good now we know d and q d = 23 + n = 23 - 9 q = 23 - n = 23 + 9

OpenStudy (anonymous):

So what do we do now?

OpenStudy (amistre64):

^^^^ look at the post above ..

OpenStudy (anonymous):

d = 23 + n = 23 - 9 q = 23 - n = 23 + 9 doesn't answer the question

OpenStudy (amistre64):

im pretty sure it does ..... what question are we trying to answer?

OpenStudy (anonymous):

How many dimes? How many quarters?

OpenStudy (amistre64):

so how many d.imes do we have?

OpenStudy (amistre64):

d = 23 + n, and n=-9

OpenStudy (anonymous):

14?

OpenStudy (amistre64):

yep how many q.uarters so we have? q = 23 - n ; and n=-9

OpenStudy (amistre64):

or if you want to use our fundamental equation: 14 + q = 46